OpenStudy (anonymous):

What is the percent of increase from 50 to 90? A. 40% B. 44% C. 56% D. 80%

OpenStudy (freckles):

percent increase is given by \[\frac{ \text{ new } -\text{ old }}{\text{ old } } \cdot 100 \text{ %}\]

OpenStudy (anonymous):

so i divide

OpenStudy (dan815):

nvm i dunno anymore I DONT KNOW LIFE

OpenStudy (anonymous):

what u mean

OpenStudy (dan815):

do what freckles says first!

OpenStudy (freckles):

no @dan815 your way was good to \[A=\frac{ \text{ new }- \text{ old }}{\text{ old }} \\ A=\frac{ \text{ new }}{\text{ old }} -1 \\ A+1=\frac{\text{new }}{\text{ old } } \\ \] \[\text{ old } (A+1)=\text{ new }\] Where your I=A+1

OpenStudy (freckles):

too!*

OpenStudy (dan815):

One more way for you 50+x% of 50 = 90 50+x/100 * 50 = 90 50*(1+x/100)=90 solve for x 2 moar way for you 50*R=90 r=90/50 r=1+k 90/50 = 50/50 + 40/50 =50/50 + 80/100 ^----- 80%! increase voila

OpenStudy (anonymous):

a 40